Circular No. 8033

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 X5 (KUDO-FUJIKAWA) J. Watanabe, National Astronomical Observatory of Japan, reports the

independent visual discovery of this comet by Shigehisa Fujikawa (observation below; report received by Watanabe prior to the issuance of IAUC 8032). T. Lovejoy, Thornlands, Qld., Australia, reports that C/2002 X5 appears faintly visible at mag ≈ 10–11 (based on the appearance of other comets in SWAN images) on SWAN images that were taken on six dates, Nov. 6–13, and posted at the SOHO website. Additional astrometry and the following (revised) preliminary parabolic orbital elements appear on MPEC 2002-Y14; the ephemeris magnitudes assume H = 7.0, n = 3.

T = 2003 Jan. 28.776 TT  ω = 188.^{o}205
Ω = 119.389 }2000.0
q = 0.18485 AU  i = 94.431
2002/03 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 12 15^{h}29.^{m}53 +46^{o}17.^{′}2 1.161 1.289 73.^{o}4 47.^{o}0 8.2
14 15 45.96 +45 20.4 1.130 1.249 72.1 48.6 8.0
16 16 02.60 +44 13.1 1.101 1.209 70.6 50.2 7.8
18 16 19.32 +42 54.5 1.075 1.168 69.0 51.8 7.7
20 16 36.01 +41 24.3 1.052 1.126 67.1 53.6 7.5
22 16 52.53 +39 42.3 1.031 1.084 65.1 55.4 7.3
24 17 08.79 +37 48.6 1.014 1.041 62.8 57.2 7.2
26 17 24.66 +35 43.5 0.999 0.998 60.4 59.0 7.0
28 17 40.08 +33 27.6 0.989 0.953 57.8 60.8 6.8
30 17 54.97 +31 01.9 0.981 0.908 55.1 62.6 6.6
Jan. 1 18 09.28 +28 27.3 0.977 0.862 52.2 64.3 6.5
3 18 22.98 +25 45.0 0.976 0.815 49.2 65.9 6.3
5 18 36.06 +22 56.0 0.979 0.767 46.0 67.3 6.1
7 18 48.52 +20 01.4 0.985 0.718 42.8 68.5 5.9
9 19 00.37 +17 01.9 0.995 0.668 39.5 69.3 5.7
11 19 11.64 +13 58.2 1.007 0.617 36.1 69.8 5.4

Total-visual-magnitude and coma-diameter estimates: Dec. 14.858 UT, 9:, 4′ (S. Fujikawa, Oonohara, Kagawa, Japan, 0.16-m reflector); 15.50, 8.5, 5′ (A. Hale, Cloudcroft, NM, 10×50 binoculars); 16.20, 7.9, 6′ (J. J. Gonzalez, Asturias, Spain, 11×80 binoculars); 16.58, 7.8, — (S. J. O’Meara, Volcano, HI, 0.10-m refractor); 16.66, 7.6, 5′.5 (M. Linnolt, Honolulu, HI, 0.11-m reflector).

2002 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green